engineering economics

A contractor has been awarded the contract to construct a 6-mile-long tunnel in the mountains. During the 5-year construction period, the contractor will need water from nearby stream. She will construct a pipeline to carry the water to the main construction yard, an analysis of costs for various pipe sizes is as follows 3. Pipe Sizes (in) 24,000 28,000 30,000 Installed cost of pipeline and pump 20,000 Cost per hour for 1.5 0.75 0.55 0.35 Salvage value of pump MARR . 10% zero Pump will operate 1500 hours Select the alternative with the least present worth of cost: (20 pts) a. 6 with net present worth of cost KD 31,990 b. 3 with net present worth of cost KD 28,265 c. 2 with net present worth of cost - KD 45,530 d. 1 with net present worth of cost - KD 13,265
Answer #1

Pipeline will operate for 1,500 hours each yeary for 5 years. So it will operate 300 hours each year.

For all four options every years pumping cost will be:

1. 1500 *1.5 = 2,250

2. 1500*0.75 = 1,125

3. 1500* .55= 825

4. 1500* .35= 525

With MARR being 10% each year's pumping cost will be discounted by 10%.

Adding 5 year discount factor: 1/(1.10)^1 + 1/(1.1)^2 + 1/(1.1)^3 + 1/(1.1)^4 + 1/(1.1)^5 = 3.7907868

So present value of total cost for each pipeline is as follows:

yearly pumping cost 5 year pumping cost after discounting installation cost Total cost

2250 8529.270231 20000 28529.27023

1125 4264.635116 24000 28264.63512

825 3127.399085 28000 31127.39908

525 1990.163054 30000 31990.16305

Based on these options second option that is 3 inch pipe installation will be cheapest. Total cost will be 28,264.64 or 28,265 to match b option

Thanks